OpenStudy (anonymous):

Solve the equation. Be sure to show all of your work. 4r + 2 = 26

OpenStudy (anonymous):

4r=26-2 r= 24/4 r= 6

OpenStudy (anonymous):

4r=26-2 4r = 24 divide by 2 both side r=6

OpenStudy (anonymous):

4r+2=26 -2 -2 4r=24 divided by 4 r=6

OpenStudy (anonymous):

yah by 4

OpenStudy (anonymous):

thanks guys so much!

OpenStudy (anonymous):

u r we com